OpenStudy (anonymous):

Two cards are drawn from a standard deck of cards. Find each probability. 1. P(both jacks or both face cards) 2. P(both either black or an ace)

OpenStudy (anonymous):

both jacks is a subset of both facecards, so the first one means P(both face cards) there are 12 face cards in a deck, so probability you get two of them is \[\frac{12}{52}\times \frac{11}{51}\]

OpenStudy (anonymous):

second is \[\frac{28}{52}\times \frac{27}{51}\]

OpenStudy (anonymous):

there are 26 black cards and two aces that are not black. that explains the 28

OpenStudy (anonymous):

the key ; 1) 11/221 2) 63/221

OpenStudy (anonymous):

got a calculator?

OpenStudy (anonymous):

ok, your answer is correct , ty

OpenStudy (anonymous):

They are equivalent.